OpenStudy (anonymous):

f(x) = -2x^3 - 5x^2 - 6x + 4 ? What are the zeros of this equation

OpenStudy (anonymous):

first by trail nd error find a zero...

OpenStudy (lgbasallote):

ooohhh sir @eliassaab in action...this is very valuable *_* *popcorn*

OpenStudy (anonymous):

There are two complex roots and one real root. Try graphing.

OpenStudy (lgbasallote):

is there a way to know that? like a formula??

OpenStudy (anonymous):

no formula first find one zero by trail nd error

OpenStudy (anonymous):

There is a general formula for cubics equations, see http://en.wikipedia.org/wiki/Cubic_function

OpenStudy (anonymous):

I used the Rational Root Theorem and Descartes' Rule of Signs. But I can't find that one positive zero.

OpenStudy (anonymous):

datz y i said try de ec way....first trail nd error den take de factor....den divide u'll get a quadratic polynomial, den split de middle term to find de odr 2 zeroes

OpenStudy (anonymous):

Wait I got it thanks eliassab

OpenStudy (anonymous):

Here it is \[x=\frac{1}{6} \left(-5-\frac{11}{\sqrt[3]{361+6 \sqrt{3657}}}+\sqrt[3]{361+6 \sqrt{3657}}\right) \]

OpenStudy (anonymous):

\[ x\approx 0.458932 \]